OpenStudy (mtalhahassan2):

Adding and subtracting Rational Expression m divide by 4 + 3 m divide by 5

OpenStudy (mtalhahassan2):

Is it would be 4m divide by 20?

OpenStudy (aum):

Is the question \[\Large \frac{m}{\frac{4+3m}{5}} \]?

OpenStudy (mtalhahassan2):

no

OpenStudy (aum):

The question is not clear.

OpenStudy (aum):

\[\frac m4 + \frac {3m}{5} \]?

OpenStudy (mtalhahassan2):

yes this one is corrct

OpenStudy (aum):

\[\frac ab + \frac cd = \frac{ad + bc}{bd}\]

OpenStudy (aum):

\[ \frac m4 + \frac {3m}{5} = \frac{5*m + 4*3m}{4*5} = ? \]

OpenStudy (mtalhahassan2):

27 m/20

OpenStudy (aum):

numerator = 5m + 12m = 17m

OpenStudy (mtalhahassan2):

thanks man

OpenStudy (aum):

\[ \frac m4 + \frac {3m}{5} = \frac{5*m + 4*3m}{4*5} = \frac{5m+12m}{20} = \frac{17m}{20} \]